Earth has a mass of 5.97×1024 kg and a radius of 6,370 km. What is its moment of

inertia if it is assumed to be a solid sphere of uniform density?


1
Expert's answer
2021-01-07T10:56:29-0500

The moment of inertia of a uniform solid sphere is:


"I = \dfrac25mr^2"

where "m = 5.97\times 10^{24}kg" is its mass, and "r = 6.37\times 10^6m" is its radius. Thus, obtain:


"I = \dfrac25\cdot 5.97\times 10^{24}\cdot (6.37\times 10^6)^2 \approx 9.7\times10^{36} kg\cdot m^2"

Answer. "9.7\times10^{36} kg\cdot m^2".